OpenStudy (brelove13):

. The half-life of a radioactive substance is four days. (a) Find the decay constant. (b) How much of an initial 2-g sample will remain after 10 days?

OpenStudy (mrcoolguy):

What do you think? Also, is that all of your question?

OpenStudy (brelove13):

yes this is all of the question and i don't even understand what it's asking

OpenStudy (mrcoolguy):

Me either. sorry man

OpenStudy (brelove13):

it's ok..my professor never never teaches what he tests us on

OpenStudy (mrcoolguy):

ikr, maybe @KendrickLamar2014 or @ShadowLegendX can help you.

OpenStudy (kendricklamar2014):

I'm not sure about this .-.

OpenStudy (brelove13):

ok thanks guys

OpenStudy (wolf1728):

Here's a good explanation of half life http://www.1728.org/halflife.htm and here's a solution: half life = 4 days How much of 2 gram sample is left after 10 days? Ending Amount = Beginning Amount / (2^n) where n = number of half-lives if half life = 4 days, after 10 days, 2.5 half lives have passed. Ending Amount = 2 grams / (2^2.5) Ending Amount = 2 grams / 5.657 Ending Amount = 0.3535 grams ************************************************************************** The decay constant (represented by the Greek letter lambda, which you don't need now) equals (natural log of 2) / half-life decay constant = 0.6931471806 / 4 More about the decay constant here: https://en.wikipedia.org/wiki/Half-life